OpenStudy (anonymous):

\[(x+xy^2)dx + e^{x^2}ydy =0\]

OpenStudy (anonymous):

\[(x+xy^2)dx + e^{x^2}ydy=0\]\[x(1+y^2)dx =- e^{x^2}ydy\]\[-\frac{x}{e^{x^2}}dx=\frac{y}{1+y^2}dy\]\[-\frac{1}{2}e^{-x^2}d(x^2)=\frac{1}{2(1+y^2)}d(y^2)\]\[\frac{1}{2}e^{-x^2}+C=\frac{1}{2}\ln |1+y^2|\]\[e^{-x^2}+C=\ln |1+y^2|\]Does it look fine?

hartnn (hartnn):

it looks correct to me.

OpenStudy (anonymous):

i dont see anything wrong here

OpenStudy (anonymous):

Oh!! Thanks!! Should I further simplify it and make y as the subject or just leave it here?

hartnn (hartnn):

no, point.... that will just complicate the equation. exponential of exponential.....

OpenStudy (anonymous):

Okay, thanks! :)

hartnn (hartnn):

welcome ^_^